OpenStudy (anonymous):

Kathy is financing $218,675 to purchase a house. She obtained a 30/10 balloon mortgage at 3.35%. What will her balloon payment be? $174,042.89 $152,249.65 $168,871.56 $168,401.44 @amistre64

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

omg are you taking the final right now?!

OpenStudy (anonymous):

adv alg with financial application?

OpenStudy (amistre64):

ok, so tell me what your understanding of a balloon payment is

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

the balloon payment has a lower price at first but after the end a large sum is owed from the loan

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

this question is about a 30/10 so its a 30 year mortgage

OpenStudy (amistre64):

balloon payment is just a payment, its the final payment we need to make.

OpenStudy (amistre64):

we make 10 years of monthly payments calculated on a 30 year loan whats our normal monthly payments on a 30 year loan?

OpenStudy (anonymous):

for a normla 30 year mortgage it would be 607.430

OpenStudy (amistre64):

thats not what i get for a monthly payment i have a formula that I developed \[B_n=Bk^n-P\frac{k^n-1}{k-1}\] for n compounding periods (30*12) = 360 we want the loan paid off, B{360} = 0 \[0=Bk^{360}-P\frac{k^{360}-1}{k-1}\] we can solve for P \[Bk^{360}\frac{k-1}{k^{360}-1}=P\] k is our compounding stuff, k = 1+.0335/12 and B is our initial loan amount

OpenStudy (anonymous):

i got 360 i was spreading that over the loan amount for individual payments... what did you get?

OpenStudy (amistre64):

P = 963.73

OpenStudy (amistre64):

yeah, your monthly payment approach is invalid. we have to deal with interest so dividing the loan amount by 360 is not appropriate

OpenStudy (amistre64):

now what we need, is the balance on the loan after we make 10*12 - 1 payments 120 payment make up all of the payments, 119 are regular monthly payments, and the last one is a balloon, it is the compounded balance after that 119th payment was sent in.

OpenStudy (anonymous):

oh, ok, so how do we figure that

OpenStudy (amistre64):

i gave you how I figure out a monthly payment with a formula i derived for it. now the ballon uses the same formula, but with the relevant information \[B_{119}=Bk^{119}-P\frac{k-1}{k^{119}-1}\] \[Balloon=B_{119}k\color{red}{\implies} Bk^{120}-Pk\frac{k-1}{k^{119}-1}\]

OpenStudy (amistre64):

you need to know how to find a monthly payment, im concerned that you dont know that yet.

OpenStudy (amistre64):

ugh, i got my geometric sum on my ballon payment flipped over. k^n -1 should be on top \[B_{119}=Bk^{119}-P\frac{k^{119}-1}{k-1}\] balloon is still B{119} * k of course

OpenStudy (anonymous):

so i use the formula you gave me with the info im given from my question then

OpenStudy (amistre64):

thats how i work it yes

OpenStudy (anonymous):

do i change b?

OpenStudy (amistre64):

how your material works it is another story :)

OpenStudy (anonymous):

in this formula what is K?

OpenStudy (amistre64):

B defines your loan amount. Bn defines the balance onthe loan

OpenStudy (amistre64):

and as stated earlier, k is the compounding rate ... re-read the posts

OpenStudy (anonymous):

right sorry

OpenStudy (anonymous):

so here is what i think i have P=963.73 B=218,675 BN=3.35% K=10 is any of the right

OpenStudy (amistre64):

P and B are right :) k = 1 + .0335/12 Bn is just a balance after n = (10*12-1) periods

OpenStudy (amistre64):

now the issue i have is that my solution is not an option. so im going to double check this with an online calculation

OpenStudy (anonymous):

ok i can see how that works

OpenStudy (amistre64):

119th monthly payment is 218675k^(10*12-1)-963.73(k^(10*12-1)-1)/(k-1) = 168 893. 73 which is only 3 cents off of the online calculator ----------------------------------- my balloon is 169 365. 22 which is just a few cents off of the online as well none of your options fit this.

OpenStudy (anonymous):

i can see... ill try my teacher